How many turns per linear meter should have a solenoid so that when placed around a vacuum cathode ray tube can maintain a 100eV electron circulating in the tube axis in a circle of 1cm radius. The solenoid carries a current of 1 A.

Ans: Aprox, 2700 turns/meter
Answer #1

Speed of the electron ..1

Here M=9.11*10-31 kg , 100eV=100*1.6*10-19 =16*10-18 J

Put these value in equation 1

9.11*10-31V2 =2*16*10-18

V2 =320*10-19/9.11*10-31

V2=35.13*1012

V=5.927*106 m/s

now magnetic force on the electron

=qVB=MV2/R

B=MV/Rq ..2

for solenoid ..3

From equation 1 and 3

N=MV/IRq

4*10-7 , R=1cm=0.01 m ,I=1.0A q=1.6*10-19C

N=9.11*10-31*5.927*106/4*10-7*1*0.01*1.6*10-19

N=2.686*103

N=~2700 turns